# Latent Beam Diffusion Models for Decoding Image Sequences

### v2

@@ B.1 Metrics and Annotations
Table 5: Comparison of CLIP-I, DINO-I and CLIP-T across different methods with the standard deviation
Method CLIP-I DINO-I CLIP-T
GILL 0.55 $\pm$ 0.30 0.49 $\pm$ 0.26 0.43 $\pm$ 0.08
StackedDiffusion 0.76 $\pm$ 0.12 0.46 $\pm$ 0.14 0.47 $\pm$ 0.09
Image Sequence Decoding
Greedy 0.81 $\pm$ 0.13 0.54 $\pm$ 0.14 0.48 $\pm$ 0.07
Nucleus 0.77 $\pm$ 0.15 0.53 $\pm$ 0.15 0.48 $\pm$ 0.08
BeamDiffusion 0.83 $\pm$ 0.14 0.62 $\pm$ 0.18 0.48 $\pm$ 0.09
To complement average scores, we analyze the standard deviation ($\sigma$) of these metrics (Table 5). A low $\sigma$ reflects stable, consistent performance between diverse sequences, while a higher $\sigma$ implies variability in quality or adaptation to task complexity.
GILL exhibits the highest $\sigma$ in CLIP-I and DINO-I, indicating its semantic similarity varies significantly between tasks, likely due to inconsistent adherence to step instructions. StackedDiffusion’s low $\sigma$ indicates its consistent performance across different tasks. BeamDiffusion strikes a balance with moderate $\sigma$ values, indicating a balance between consistency and variation across tasks.
It is worth noting that CLIP-T generally shows lower standard deviation across models, since this metric is explicitly minimized during the diffusion model training. As a result, CLIP-T scores tend to be more stable regardless of the task-specific variation.
Together, the mean and standard deviation metrics suggest BeamDiffusion offers the best combination of high semantic fidelity, text alignment, and stable adaptation across different sequences, making it the most robust method for sequential image generation.
@@
Gemini evaluation.
Gemini’s selection process for annotating the best beam search configuration and the best generation method closely mirrored that of human annotators, with one key adjustment: Gemini tends to default to "both good" or "both bad" when such options are available. To ensure more decisive and accurate results, we removed these options from the prompt. We also refined our comparison methodology. Initially, presenting all five sequences at once led to biased evaluations, as the model disproportionately favored the last two sequences. To mitigate this, we adopted a pairwise evaluation strategy, comparing two sequences at a time. This approach resulted in more balanced and reliable assessments across all comparisons.
Figure 7 shows the prompt used in Gemini to guide the evaluation of image sequences based on visual and semantic consistency.
Gemini Prompt: "Evaluate two image sequences using the following criteria: Visual Consistency: Objects, ingredients, backgrounds, colors and styles should remain consistent across frames. Frames should transition smoothly, resembling a video-like progression with no sudden changes or missing elements. Semantic Consistency: Images must logically follow the described steps, clearly and accurately. Steps of the process: {steps} After evaluating all sequences, select the one with the strongest overall consistency (visual and semantic). Pay close attention to even minimal differences, as the sequences are very similar. If both sequences are equally good or bad, choose the one that is slightly better. Choose exactly one of the following options: - "First sequence" - "Second sequence" "
Figure 7: Gemini prompt used to select the best beam search configuration and best method.
B.2 Datasets
Recipes and DIYs.
Each manual task in both the Recipes and DIYs datasets includes a title, a description, a list of ingredients, resources and tools, and a sequence of step-by-step instructions, which may or may not be illustrated. The Recipes dataset comprises approximately 1,400 tasks, with an average of 4.9 steps per task, totaling 6,860 individual steps. Most tasks feature an image for each step. The DIYs dataset includes a variety of tasks, from simpler projects to more complex ones, with detailed steps.
@@ B.2 Datasets
Appendix C Impact of Latent Selection on Beam Search Performance
C.1 Impact of Latent Groups on Performance
Annotations.
Following the initial analysis of beam search configurations (Sec. 5.3), where we used latent indexes 0, 1, 2, and 3, we further investigated the impact of using different latent configurations on the generated sequences. While latents 0, 1, 2, and 3 show promising results, latents 1, 3, 5, and 7 exhibit significantly poorer performance. In particular, beam search configurations for these underperforming latents result in very limited selection rates, indicating they are rarely chosen in both human and Gemini annotations. For human annotations across 6 tasks, the only configuration that yielded any selection was the one with 4 steps back and a beam width of 3, achieving a selection rate of 8.33%. All other configurations resulted in 0% selection. In contrast, none of the configurations showed any non-zero selection rate under Gemini annotations, reinforcing the lack of preference for these latents.
Figure 8: Impact of using 1, 3, 5, and 7 latent configurations on image generation.
Similarly, for Gemini annotations evaluated on the same 6 sequences, the configurations do not yield any notable results, with no selection rates. When the evaluation is extended to 35 sequences, the performance for latents 1, 3, 5, and 7 remains sparse, with a single configuration (2 steps back, beam width of 3) being selected, with a selection rate of 2.86%. Figure 8 visually demonstrates the poor results, where the images appear nearly identical, further supporting the findings of limited diversity and low performance, which aligns with the findings of [33, 2].
Automatic Metrics.
Given the observed negative impact of higher latents on performance in human annotations, we investigated how latents influence text-image alignment using the DAAM (Diffusion Attention Attribution Map) framework [46]. DAAM interprets text-to-image diffusion models by analyzing cross-attention maps during the denoising process, generating heat maps that reveal how strongly each prompt token affects different regions of the output image.
By applying DAAM, we examined whether higher latents distort or weaken the influence of key tokens during generation. The resulting heat maps allowed us to visually assess token-level contributions, helping us understand whether the model continues to attend to the most semantically important parts of the prompt under varying latent configurations..
To quantify text alignment, we calculated the average heat map maximum intensity for all tokens, providing a numerical measure of how well the model preserved the key textual elements in the generated image. Then, for each sequence of generated images, we computed the average alignment score across all images in the sequence. While this serves as an objective metric for evaluating text adherence, it remains a partial measure, as it focuses solely on text alignment without accounting for other important aspects of image quality, such as coherence, realism, or artistic fidelity. Figure 9 shows the average impact of different latent groups on text alignment. The results reveal a clear pattern, latents from earlier iterations contribute more effectively to aligning the text with the generated image, while text alignment weakens as we move to higher latent ranges. Specifically, latents in the 0–3 range showed the strongest alignment, with an average score of 1.544, followed by the 4–7 range (1.523). In contrast, latents from the highest index range (16–19) had the weakest influence on alignment, with a significant drop in average scores. This suggests that earlier-stage latents play a crucial role in preserving text fidelity, whereas later-stage latents may dilute or distort the contribution of key tokens.
Figure 9: The average influence of latent indices on text alignment across task data.
In Figure 13, we analyze text alignment across multiple steps using heat maps corresponding to different latent groups, revealing how the connection between the input text and the generated image evolves throughout the process. Certain areas, such as the background and the bowl, show little to no heat map activation, indicating that these regions are less influenced by the text and more shaped by the model’s learned priors and overall scene composition. Since the model generates new images by building on latents from previous steps, these early decisions about text relevance persist throughout the process, influencing later refinements. Additionally, when examining generations using later latent groups, we find that this information remains more stable across all steps. This suggests that latents at these stages carry more structural and stylistic details, focusing on consistency rather than continuous adjustments based on the input text.
While later latents are effective at preserving visual consistency from previous steps, they are less capable of incorporating new details from the input prompts. For example, later latent groups (16-19) maintain stability in visual elements throughout the process, preserving previously established features. However, this comes at the expense of adapting to new or changing textual instructions. In contrast, early latent groups (0-3) demonstrate strong alignment between the text and the generated image, with consistently high heat map intensity across all steps. This highlights their role in grounding the image in the input text. For later latent groups (4-7), however, we observe a drop in heat map intensity, particularly by step 3, suggesting the model shifts focus away from strict textual adherence and toward refining finer visual details.
C.2 Latent Selection on Image Sequence Generation
Figure 11: Heat maps illustrating cross-attention in latent diffusion across different decoding methods. The visualizations show variations in how each approach preserves or redistributes attention to textual cues throughout the generation process.
Figure 11 illustrates how different latent search methods affect text-image alignment. The heat maps show how attention patterns shift depending on the decoding strategy, revealing variations in how well each method preserves focus on key textual cues. Notably, in step 3, BeamDiffusion exhibits superior alignment, maintaining a stronger connection to the prompt and ensuring greater sequence coherence compared to other methods.
While Beam Diffusion generally achieves better visual consistency and stronger text adherence in earlier steps, it slightly underperforms in step 5, failing to capture the "fireworks" mentioned in the prompt. This highlights how even methods with strong overall performance can exhibit localized weaknesses, underscoring the importance of analyzing alignment step by step.
Overall, these findings highlight the significance of latent selection strategies in optimizing image generation. The results indicate that prioritizing the most significant latents could enhance the coherence of visual sequences and improve text alignment. These observations underline the influence of latent search strategies on the semantic accuracy and coherence of generated outputs.
C.3 Effect of Cumulative Latent Access
As shown in Figure 12, limiting the model to latents from only the first step results in weak text-image alignment and incoherent visual transitions. Gradually expanding the latent set to include more steps improves sequence quality, particularly when moving from two to three steps. However, we observe that including latents from all previous steps does not yield further improvements in coherence, suggesting diminishing returns with increasing temporal access.
In contrast, BeamDiffusion uses latents from only the two most recent steps and achieves results that are comparable, or even superior, to the full cumulative case. This indicates that a narrow temporal window is sufficient to maintain strong alignment and visual consistency. By avoiding the computational cost of considering all previous steps, BeamDiffusion finds a favorable trade-off between quality and efficiency.
Overall, this experiment highlights that cumulative latent access enhances generation quality up to a point, but full historical memory is unnecessary. Strategies such as BeamDiffusion, which take advantage of a limited but informative temporal context, can produce high-quality sequences while significantly reducing computational overhead. These findings also underscore the importance of the non-linear sequence denoising strategy introduced in Section 5.4, which enables more targeted and effective reuse of past latents without the need for full sequence access.
Appendix D Minimizing Decoding Risk
Greedy Decoding and Nucleus Sampling.
BeamDiffusion.
Random Seeds for mid-sequence Generation.
Appendix E Broader Impacts and Challenges
@@
E.1 Impacts
E.2 Contextual Scene Descriptions
One of the challenges encountered concerns the generation of captions that accurately describe the steps involved in the sequence. Often, a single step may consist of multiple sub-steps, each contributing to the overall generation. This makes it difficult to write a single concise caption that captures the full scope of each step. The challenge lies in ensuring that the caption not only reflects the specific action or transformation occurring at each step but also conveys the progression and complexity of the entire sequence. Additionally, if we divide the steps into smaller sub-steps to better describe the process, the sequence becomes very extensive, making it even harder to maintain clarity and readability while ensuring comprehensive coverage of the entire generation process.
To address this challenge, we leverage Gemini [34], specifically the Gemini 1.5 Flash model, to refine step descriptions into visually detailed prompts while considering sequence context. Specifically, Gemini produces a contextualized caption $c_{j}$ based on the current step description $s_{j}$ and all preceding steps as follows:
$\small c_{j}=\phi(s_{j}|\{s_{1},s_{2},\dots,s_{j-1}\})$ (7)
E.3 Domains.
@@ E.2 Contextual Scene Descriptions
Appendix F Application Examples
@@
This section presents additional visual examples and qualitative results that complement the main paper. Images and other visual aids are essential tools that significantly improve comprehension across a variety of domains. Whether breaking down complex tasks, illustrating a sequence of events, or clarifying intricate concepts, visual representations provide clear, step-by-step guidance that enhances understanding [24]. From technical procedures and instructional design to storytelling and educational content, multi-scene image sequences help readers grasp information more intuitively, transforming even the most complex or dense topics into something more accessible. However, generating a sequence of images that not only aligns with each textual instruction but also maintains overall coherence remains a major challenge [23]. In this appendix, we include expanded examples to further illustrate how our model addresses this issue. These include additional visualizations of the beam search process and extended qualitative comparisons across domains such as recipes, visual storytelling, and DIY tasks. Together, these examples highlight the strengths and limitations of current approaches in achieving textual alignment and visual consistency in sequential image generation.
Figure 13: Visualization of cross-attention in latent diffusion across different sets of latents. The heat maps reveal how attention patterns evolve at various sequence steps, highlighting the alignment between text prompts and distinct latent representations.
F.1 Beam Search Tree
(a) The beam denoising search tree.
In Figure 14(a), we provide another example of how the latent search occurs within the beam denoising process. This illustrates the BeamDiffusion model at work, showing the denoising of the initial seed to latents and how different beam paths are decoded. The figure highlights how our method explores the latent space, iterating through various latents based on the beam search strategy.
In Figure 14(b), we observe that the generation of the selected image at step 4 behaves similarly to an isolated diffusion process. However, it is important to note that this process is influenced by multiple prompts. Specifically, the image generation at step 4 is conditioned on the prompts from steps 1, 3, and 4. This demonstrates how the latent search process incorporates context from various steps, showing that even though each generation step seems independent, it is shaped by previous steps, enabling a more coherent and consistent generation of the final image.
F.2 Sequence Examples
@@
To provide a comprehensive comparison between all models, we present several examples across different domains in Figures 18-26. These figures show case how each method performs in terms of maintaining sequence coherence throughout sequential image generation.
In our overall qualitative analysis, we observe that both Greedy/CoSeDlen=1 and Nucleus Sampling methods are generally good at following the textual input across steps, as shown in Figures 18, 18, 22, and 22. They align well with the step instructions but often struggle with visual consistency, as the generated images show noticeable shifts or inconsistencies between steps. This lack of visual cohesion detracts from the overall coherence of the sequence, even though they adhere reasonably well to the textual prompt. In contrast, GILL and StackedDiffusion exhibit a different behavior. While both maintain high visual consistency, generating sequences in which the images are visually similar, they struggle to accurately follow the step-by-step textual instructions. GILL, in particular, tends to produce images that lack the necessary variability to incorporate new information from each step, resulting in sequences that are visually cohesive but contextually incorrect or incomplete. StackedDiffusion performs slightly better than GILL in this regard, capturing some of the step-wise changes, but still often fails to fully reflect the evolving textual input across the sequence.
Finally, BeamDiffusion strikes a strong balance between text adherence and visual consistency. As shown in Figures 18, 18, 22, and 22, it generally produces the most coherent sequences compared to other methods. In most cases, BeamDiffusion effectively follows the textual instructions while maintaining visual consistency across the steps. This combination of textual alignment and visual stability leads to more coherent and contextually accurate image sequences. While not without its limitations, BeamDiffusion demonstrates a more reliable progression throughout the image generation process than the other approaches. However, it is important to note that all methods face challenges in maintaining coherence within the DIY domain, as seen in Figures 26, 26, and 26. The complexity and diversity of DIY tasks present unique difficulties for the models, making it harder to both accurately follow textual instructions and ensure consistency throughout the sequence. While the models perform well in other domains, the intricacies of the DIY domain result in some inconsistencies, particularly when capturing step changes.
Figure 15: Example of the Recipes domain task "Vegan Basil Pesto and Sun-Dried Tomato Pinwheels", with a sequence of 4 steps.
Figure 16: Example of the Recipes domain task "Tangy lemon sheet cake", with a sequence of 4 steps.
Figure 17: Example of the Recipes domain task "Zucchini, Pecorino & Mint Salad", with a sequence of 4 steps.
@@
Figure 18: Example of the Recipes domain task "Bone Broth", with a sequence of 3 steps.
Figure 19: Example of the out-of-domain VIST story "Klamath Project Construction", with a sequence of 5 steps.
Figure 20: Example of the out-of-domain VIST story "Yosemite Weekend", with a sequence of 5 steps.
Figure 21: Example of the out-of-domain VIST story "Fireworks - 4th July in Washington, DC", with a sequence of 5 steps.
Figure 22: Example of the out-of-domain VIST story "Independence Day Parade", with a sequence of 5 steps.
Figure 23: Example of the out-of-domain DIY task "How to Clean Silk Rugs", with a sequence of 2 steps.
@@
Figure 24: Example of the out-of-domain DIY task "How to Polish Wood", with a sequence of 3 steps.
@@
Figure 25: Example of the out-of-domain DIY task "How to Preserve Basil", with a sequence of 6 steps.
@@
Figure 26: Example of the out-of-domain DIY task "How to Clean Natural Stone", with a sequence of 5 steps.

### v3

@@ Appendix E Metrics and Annotations
Even after applying this clipping, we observe that automatic metrics remain imperfectly aligned with human and LLM-as-a-judge evaluations, underscoring the ongoing challenge of developing reliable metrics for evaluating sequences of images.
@@
LLM-as-a-judge Evaluation.
The LLM-as-a-judge selection process for annotating the best beam search configuration and generation method closely mirrored that of human annotators, with one key adjustment: the model tended to default to "both good" or "both bad" when such options were available. To encourage more decisive outcomes, we removed these options from the prompt. We also refined our comparison methodology. Presenting all five sequences simultaneously initially led to biased evaluations, as the model disproportionately favored the last two. To address this, we adopted a pairwise evaluation strategy, comparing two sequences at a time. This adjustment yielded more balanced and reliable assessments across all comparisons.
Figure 13 presents the LLM-as-a-judge prompt for evaluating image sequences on visual and semantic consistency.
LLM-as-a-judge Prompt:
Evaluate two image sequences using the following criteria:
Visual Consistency: Objects, ingredients, backgrounds, colors, and styles should remain consistent across frames. Frames should transition smoothly, resembling a video-like progression with no sudden changes or missing elements. Semantic Consistency: Images must logically follow the described steps, clearly and accurately. Steps of the process: {steps} After evaluating both sequences, select the one with the strongest overall consistency (visual and semantic). Pay close attention to even minimal differences, as the sequences are very similar. If both sequences are equally good or bad, choose the one that is slightly better. Choose exactly one of the following options:
  - "First sequence"
  - "Second sequence"
Figure 13: LLM-as-a-judge prompt used to select the best beam search configuration and best method.
Appendix F Datasets
Recipes.
Each manual task in the Recipes datasets includes a title, a description, a list of ingredients, resources and tools, and a sequence of step-by-step instructions, which may or may not be illustrated. The Recipes dataset comprises approximately 1,400 tasks, with an average of 4.9 steps per task, totaling 6,860 individual steps. Most tasks feature an image for each step.
@@ Appendix F Datasets
Appendix G Challenges
@@
Captions.
One of the challenges encountered in this process is generating captions that accurately describe the steps involved in the sequence. Often, a single step may consist of multiple sub-steps, each contributing to the overall generation. This makes it difficult to write a single concise caption that captures the full scope of each step. The challenge lies in ensuring that the caption not only reflects the specific action or transformation occurring at each step but also conveys the progression and complexity of the entire sequence. Additionally, if we divide the steps into smaller sub-steps to better describe the process, the sequence becomes very extensive, making it even harder to maintain clarity and readability while ensuring comprehensive coverage of the entire generation process.
Domains.
@@
(a) Example 1: 50-step generations exhibit blur propagation, while 100 steps reduce blur and sharpen frames. 150 steps show no substantial gain over 100.
(b) Example 2: 50 steps produce blurry, inconsistent frames, while 100 steps improve fidelity and coherence. 150 steps do not yield further noticeable improvements.
Figure 14: Qualitative comparison of 50 vs. 100 denoising steps across two examples. Increasing steps to 150 does not produce substantial gains over 100, indicating diminishing returns.
Appendix H Implementation Details
Base Models.
Our method is primarily implemented on top of FLUX.1-dev[2], a diffusion transformer–based model released by Black Forest Labs. We extend its open-source implementation to support beam-based generation, latent trajectory tracking, and contrastive scoring mechanisms as described in the main text. Unless otherwise specified, we follow the default configuration of FLUX.1-dev with 100 denoising steps, chosen to balance generation quality and computational efficiency. While the default 50 steps generally produce globally coherent images, we observe that finer structures often remain underdeveloped, resulting in blur and reduced visual consistency across sequences. Increasing to 100 steps substantially sharpens details and improves sequence coherence, as illustrated in Figures 14(a) and 14(b). Further increasing to 150 steps provides no additional gains, confirming that 100 steps is a reasonable trade-off. To ensure that the choice of decoding method is not dependent on a specific backbone, we also implement BeamDiffusion on Stable Diffusion v2.1[39], a widely used U-Net–based latent diffusion model, using its standard parameters. This setup verifies that our conclusions regarding decoding strategies generalize across different generative architectures.
Latent Trajectory Reuse.
Classifier Training.
To train the contextual classifier, we employ a sequence-aware training approach that enables the model to capture dependencies between steps in the generation process. The classifier $\varphi$ is primarily trained on the Recipes dataset, which provides rich sequential image-text pairs of cooking steps, allowing it to learn meaningful context transitions in a structured procedural domain. Additionally, a small amount of data from the VIST dataset is included during training to introduce limited variability in narrative structure. The classifier takes as input a candidate image $I_{l}$, its textual description $s_{l}$, and the previous sequence images and respective step text ($B_{l-1}^{(b)}$). It then scores $I_{l}$ based on how well it fits within the evolving trajectory.
During training, we use labels $l_{d,l}$ indicating whether $I_{l}$ is the correct continuation in the sequence for task $d$ and step $l$. The model is trained to minimize the cross-entropy loss:
Hardware.
All experiments were conducted on a NVIDIA A100 GPU with 40GB of memory.
Appendix I Broader Impacts
Appendix J Application Examples
@@ Appendix I Broader Impacts
This section presents additional visual examples and qualitative results that complement the main paper. Images and other visual aids are essential tools that significantly improve comprehension across a variety of domains. Whether breaking down complex tasks, illustrating a sequence of events, or clarifying intricate concepts, visual representations provide clear, step-by-step guidance that enhances understanding [27]. From technical procedures and instructional design to storytelling and educational content, multi-scene image sequences help readers grasp information more intuitively, transforming even the most complex or dense topics into something more accessible. However, generating a sequence of images that not only aligns with each textual instruction but also maintains overall coherence remains a major challenge [26]. In this appendix, we include expanded examples to further illustrate how our model addresses this issue. These include additional visualizations of the beam search process and extended qualitative comparisons across domains such as recipes and visual storytelling. Together, these examples highlight the strengths and limitations of current approaches in achieving textual alignment and visual consistency in sequential image generation.
J.1 Sequence Examples
@@
To provide a comprehensive comparison between all models, we present several examples across different domains in Figures 16-18, highlighting how each method maintains sequence coherence in sequential image generation.
For the Recipes domain ("Buttermilk chicken wings", Figure 16), most models incorrectly depict the chicken as already cooked in step 2, instead of starting raw and gradually appearing cooked. BeamDiffusion correctly captures this progression, faithfully reflecting the cooking process across all steps. Similarly, in the Recipes task "Sweet potato and orange soup" (Figure 16), GILL, StackedDiffusion, 1P1S, and StoryDiffusion misrepresent the ingredient being cut, while BeamDiffusion consistently depicts a person cutting a sweet potato and maintains stepwise visual consistency without hallucinations.
For the VIST stories, BeamDiffusion demonstrates strong identity and style preservation. In "4th of July 2006" (Figure 18), it is the only method to maintain the same child between step 2 and step 4. In "Yosemite Weekend" (Figure 18), both BeamDiffusion variants maintain coherent tree styles across steps, whereas other models show a noticeable break between step 1 and step 2.
Overall, while baseline methods may partially adhere to textual instructions or maintain visual similarity, BeamDiffusion consistently balances stepwise textual alignment and visual coherence, producing sequences that are both contextually accurate and visually stabl
@@
Figure 15: Example of the Recipes domain task "Buttermilk chicken wings", with a sequence of 4 steps.
@@
Figure 16: Example of the Recipes domain task "Sweet potato and orange soup", with a sequence of 4 steps.
@@
Figure 17: Example of the VIST story domain "4th of July 2006", with a sequence of 5 steps.
@@
Figure 18: Example of the VIST story domain "Yosemite Weekend", with a sequence of 5 steps.